OpenStudy (countrygirl1431):

Please help me Algebra 2

OpenStudy (countrygirl1431):

Solve each equation. 8x2 - 24 = 3x2 - 9

OpenStudy (danjs):

put it first in this form ax^2 + bx + c = 0

OpenStudy (countrygirl1431):

i dont even get this at all

OpenStudy (danjs):

just move parts around to get it into the standard quadratic form a*x^2 + b*x + c = 0 8x^2 - 3x^2 - 24 + 9 = 0 5x^2 - 15 = 0

OpenStudy (countrygirl1431):

ok i have to find what x equals

OpenStudy (danjs):

yes, you can always use the quadratic formula to find that , do you remember this thing \[x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (countrygirl1431):

this is the first time ive ever done this lol i have no clue what im doing

OpenStudy (danjs):

an equation in form a*x^2 + b*x + c = 0 has solutions \[x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a } \]

OpenStudy (danjs):

you have 5x^2 - 15 = 0 a *x^2 + b*x + c = 0 5 *x^2 + 0*x + (-15) = 0 you can put in those values for a,b,c a=5, b=0, c=-15

OpenStudy (countrygirl1431):

ok that part makes since

OpenStudy (danjs):

there no examples in your book or notes to follow and practice?

OpenStudy (countrygirl1431):

no

OpenStudy (danjs):

here are a few, all those equations are the same form, quadratic ax^2 + bx + c = 0 solving using factoring, or the quadratic formula like we using here

OpenStudy (countrygirl1431):

that just makes it a little more confusing for me